# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v1

@@ II The model
where the distance $\mathbf{r}_{ij}$ is given by the nearest image convention using periodic boundary conditions (PBC). $\Gamma>0$ is the interaction strength, $b$ is a parameter that can be tuned to shift the spectrum of $J$, $\tilde{N}=\sum_{i}\tilde{J}_{ij}$ is Kac’s renormalization factor and $\alpha$ is the coefficient that tunes the range of interaction. The alternating sign makes the interactions antiferromagnetic when spins are separated by an odd number of lattice parameters and ferromagnetic when spins are separated by an even number of parameters, see Fig. 1. According to the classification of long-range interactions valid for both classical and quantum models, the strong long-range regime corresponds to $\alpha<d=1$ [6, 8]. This regime is characterized by the loss of additivity, although extensivity is preserved by Kac’s renormalization factor, $\tilde{N}$, ensuring a well-defined thermodynamic limit. In the limit $\alpha\to\infty$, the Hamiltonian [Eq. (1)] corresponds to the nearest-neighbor Ising chain with transverse and longitudinal fields. The tunable-range staggered interactions that we consider here are a generalization of the antiferromagnetic nearest-neighbor interactions in unfrustrated lattices. The alternating sign prevents frustration and allows for the formation of two sublattices, as sketched in Fig. 1. For vanishing fields, the ground state is the antiferromagnetic configuration, with the spins fully polarized along the $x$ axis in alternating directions for even and odd spins. Staggered tunable-range interactions have been proposed previously for the Heisenberg model [18, 39]
@@ III.1 Canonical partition function
$Z=\exp\left(-N\beta f[\bar{u}_{k}]\right)\prod_{p=0}^{M-1}\sqrt{2D_{p}}\;,$ (5)
where $f[\bar{u}_{k}]=\min_{\left\{u_{k}\right\}}f[u_{k}]$ and $f[u_{k}]=f(u_{0},\ldots,u_{M-1})$. By noticing that in the thermodynamic limit $f[\bar{u}_{k}]$ is the free energy per particle: $f[\bar{u}_{k}]=\lim_{N\to\infty}-(N\beta)^{-1}\ln Z$, it becomes apparent that computing the partition function is reduced to a minimization of the variational free energy per particle
@@ III.2 Ground State Phase Diagram
In order to study the ground state of the model, we define the variational ground-state energy as the zero-temperature limit of the variational free energy $e_{0}(u)=\lim_{\beta\to\infty}f(u)$, such that
@@
It can be shown that the global minimum, $\bar{u}$, is proportional to the staggered magnetization (See App. C)
@@
We have thus found the existence of a first-order QPT on a finite portion of the critical line. We emphasize that this is valid for all strong long-range models, $\alpha<1$. This constitutes a qualitative difference with respect to the nearest-neighbor limit, $\alpha\to\infty$, of the model, in which the QPT is second-order along the full critical line except for the point where $\omega_{z}=0$, when the model is classical [36, 35]. This is sketched in Fig. 2(b) for comparison. It is worth noting that there is some evidence that the phase transition could be mixed order [45, 46] around the classical point in the nearest-neighbor limit, but not purely first order like we have just described for the strong long-range regime [47].
@@
For $\alpha=0$ it is possible to do a classical analysis of the model that predicts the same phase diagram that we just described, with distinct finite regions of first- and second-order phase transitions. See App. D for more details. This illustrates that a classical analysis can serve as an exploratory tool for spin models with all-to-all interactions. It offers an alternative perspective on the phase transition based on the magnetizations of each sublattice rather than the order parameter (the staggered magnetization in this case). In any case, it is not a replacement for an exact solution due to the fact that its application is limited to $\alpha=0$ and that it does not allow for the computation of susceptibilities.
@@ III.3 Analysis of correlations
The susceptibility is proportional to the Kubo correlator [48][Chap. 4], and hence a measure of correlations between spins. For a translation invariant model the susceptibility can be computed analytically [49][Chap. 6] (see details in App. C). At zero temperature it can be written as
@@ IV Numerical solution for arbitrary interaction range
Having established a significant difference between the phase diagram of the strong long-range and nearest-neighbors models, it is natural to wonder how this difference is interpolated for arbitrary values of the range of interactions, $\alpha$. Since only the strong long-range regime is analytically tractable, in this section we resort to zero-temperature variational quantum Monte Carlo (qMC) [53] simulations with a visual transformer (ViT) [38, 54, 55] ansatz for finite system sizes to study the model in the full range of interactions. We show the results obtained for spins with $s=1/2$ using this architecture, given its recent success in describing spin-1/2 chains with long-range interactions [30]. Details on the particular implementation used here can be found in App. F.
@@
In Fig. 4 we observe how $\langle m_{s}^{2}\rangle$ behaves for different slices of the phase diagram (cf. Fig. 2) as a function of the different values of $\alpha$ for a fixed chain size of $N=50$. In panels (a), (b) and (c), which correspond to vertical slices (fixed transverse field), we see that for certain values of $\alpha$ the nature of the phase transition evolves from first to second order as the transverse field, $\omega_{z}$, increases. Panel (a), corresponding to $\omega_{z}/(s\Gamma)=0.2$, merits special attention. It shows that the discontinuity characteristic of first-order phase transitions is clearly present up to $\alpha=2.5$ (See App. G for further confirmation). In panels (b) and (c) we see how this transition becomes second order, with a continuous change of the order parameter at the critical point. If instead of comparing a given value of $\alpha$ across panels we focus on any given panel, (a), (b) or (c), we observe that the transition develops a discontinuity as $\alpha$ is lowered. In all cases the transition is clearly discontinuous for $\alpha\leq 1$, in agreement with the analytical predictions.
@@
To better certify the order of the different phase transitions, beyond a visual analysis of the discontinuities (or lack thereof) of the order parameter, we perform a finite-size scaling analysis. In Fig. 5 we show this analysis for $\alpha=2$ (See App. G for other values of $\alpha$). As it can be seen from panel (a), all the curves corresponding to $N=50,70,100$ coincide, showing an independence with size characteristic of first-order transitions, in addition to the discontinuity in the order parameter. In contrast, in panel (c), the curves exhibit the typical size dependence of second-order phase transitions. The continuous change of the order parameter becomes increasingly non-analytical as the size increases. The crossover point marks the second-order critical point expected in the thermodynamic limit.
@@
Figure 6 summarizes the numerical results discussed in this section, along with additional results reserved for App. G. We sketch the critical line for different values of $\alpha$, highlighting the regions where it is first and second order. We have established that the first-order phase transition at finite transverse field is present for $\alpha\leq 2.5$. The portion of the critical line that is first order decreases progressively with increasing $\alpha$ until for $\alpha\gtrsim 3$ the full critical line is second order. This is remarkable because it indicates that the first-order phase transition is present not only in the strong long-range regime but also beyond the mean-field threshold and presumably (see the next paragraph) in the full weak long-range regime. Although the precise boundaries, $\alpha_{\rm MF}$ and $\alpha^{*}$, between the regime in which the model exhibits mean-field critical exponents and the weak long-range regime and between the weak long-range and short-range regimes, have not been established for the antiferromagnetic model under consideration, it is known that for the analogous ferromagnetic model they lie at $\alpha_{\rm MF}=5/3\approx 1.66$ and $\alpha^{*}=3$ [8]. Additionally, we have witnessed the shrinking of the antiferromagnetic phase as the model approaches the short range regime, driven by a decrease of the critical transverse field at zero longitudinal field from $\omega_{z}/(s\Gamma)=2$ for $\alpha>1$ to $\omega_{z}/(s\Gamma)=0.5$ for $\alpha\gtrsim 7$.
@@ V Discussion
The fact that similar behaviour has been observed in a model with only antiferromagnetic nearest-neighbor and ferromagnetic next-nearest-neighbor interactions [31] suggests that the key ingredient that stabilizes a first-order phase transition is the presence of intrasublattice ferromagnetic interactions, in contrast with models with only intersublattice antiferromagnetic interactions. To verify this hypothesis we have also considered Hamiltonian (1) but without intrasublattice ferromagnetic interactions, i.e. with $J_{ij}=0$ for $i+j$ even. This model is not tractable with the analytical method used in Sec. III, so an exact solution in the strong long-range regime is not possible. Nevertheless, a classical analysis of the $\alpha=0$ limit, analogous to the one described in App. D, shows that the phase transition is always second order, except for the point of vanishing transverse field, $\omega_{z}=0$, where the model is classical. Thus, we can conclude that intrasublattice ferromagnetic interactions (on top of intersublattice antiferromagnetic interactions) are the key ingredient to generate a first-order phase transition. In that sense, in the tunable-range unfrustrated antiferromagnetic Ising model of Hamiltonian (1), $\alpha$ acts as a tuning knob to change the ratio between inter- and intrasublattice interactions. For low enough $\alpha$, intrasublattice ferromagnetic interactions become large enough to stabilize a first-order phase transition. The same mechanism might explain the first-order phase transitions reported in other quantum long-range models with staggered interactions [20, 21, 22].
@@ Appendix A Solving the staggered antiferromagnetic Ising chain
$Z=\int\prod_{k=0}^{M-1}\sqrt{\frac{ND_{k}}{\pi}}du_{k}\,\exp\left\{-N\beta f[u_{k}]\right\}\;,$ (18)
@@
Appendix C Analytical calculation of the magnetization and the susceptibility
@@
Appendix D Classical analysis of the phase transition for $\alpha=0$
@@
This Hamiltonian commutes with the total spin operators, $J_{\Lambda}^{2}=(J_{\Lambda}^{x})^{2}+(J_{\Lambda}^{y})^{2}+(J_{\Lambda}^{z})^{2}$. Consequently, it connects only states with the same total spins $J_{A}^{2}$ and $J_{B}^{2}$. Using exact diagonalization for finite sizes, we show in Appendix E that the ground state always lies on the subspace of maximum total spins. This implies that the ground state properties of the model are perfectly captured by a model of two spins of sizes, $j_{A}=j_{B}=sN/2$, interacting according to Eq. (39). In the thermodynamic limit, $N\to\infty$, the description is further simplified by the fact that these spins can be described exactly in their classical limit [61]. We can therefore study the energy resulting from replacing the spins with classical magnetizations $J_{\Lambda}^{\gamma}\to sN/2m_{\Lambda}^{\gamma}$. The magnetizations are unit vectors that we can assume to lie on the $x,z$ plane $\boldsymbol{m}_{\Lambda}=(\cos\theta_{\Lambda},0,\sin\theta_{\Lambda})$. With this, the energy per site reads
@@
Appendix E Verification using exact diagonalization
For $\alpha=0$ there are two equivalent descriptions of the model in terms of either individual spins at each site, as in Hamiltonian (1), or in terms of total spin operators of each sublattice, as in Hamiltonian (39). As explained in App. D, the total spin of each sublattice, $J_{\Lambda}^{2}$, is a conserved quantity and the Hamiltonian is block diagonal with blocks of fixed total spins. Here we diagonalize the subspace of maximum total spins, $j_{A}=j_{B}=sN/2$ and compare its low-energy spectrum to the low-energy spectrum of the full Hamiltonian (1). In Fig. 8 we show that the ground-state energies of the two Hamiltonians coincide, with differences appearing in the multiplicity of excited states. We show this for $N=10$ and $s=1/2$ fixing $\omega_{x}/(s\Gamma)=0.2$ and varying $\omega_{z}/(s\Gamma)$ and vice versa, but the same behavior is observed for any accessible system size, $N$, and spin size, $s$, and for any values of $\omega_{x}/(s\Gamma)$ and $\omega_{z}/(s\Gamma)$ of the phase diagram. This implies that the low energy physics are well captured by the subspace of maximum total spins, $j_{A}=j_{B}=sN/2$. This fact is exploited in App. D to provide a classical description of the ground state phase diagram.

### v2

@@ II The model
where the distance $\mathbf{r}_{ij}$ is given by the nearest image convention using periodic boundary conditions (PBC). $\Gamma>0$ is the interaction strength, $b$ is a parameter that can be tuned to shift the spectrum of $J$, $\tilde{N}=\sum_{i}\tilde{J}_{ij}$ is Kac’s renormalization factor and $\alpha$ is the coefficient that tunes the range of interaction. The alternating sign makes the interactions antiferromagnetic when spins are separated by an odd number of lattice parameters and ferromagnetic when spins are separated by an even number of parameters, see Fig. 1. According to the classification of long-range interactions valid for both classical and quantum models, the strong long-range regime corresponds to $\alpha<d=1$ [6, 8]. This regime is characterized by the loss of additivity, although extensivity is preserved by Kac’s renormalization factor, $\tilde{N}$, ensuring a well-defined thermodynamic limit. In the limit $\alpha\to\infty$, the Hamiltonian [Eq. (1)] corresponds to the nearest-neighbor Ising chain with transverse and longitudinal fields. The tunable-range staggered interactions that we consider here are a generalization of the antiferromagnetic nearest-neighbor interactions in unfrustrated lattices. The alternating sign prevents frustration and allows for the formation of two sublattices, as sketched in Fig. 1. For vanishing fields, the ground state is the antiferromagnetic configuration, with the spins fully polarized along the $x$ axis in alternating directions for even and odd spins. Staggered tunable-range interactions have been studied previously for the Heisenberg model [39, 40, 41, 42, 18, 43, 44, 45, 46, 47].
@@ III.1 Canonical partition function
$Z=A\exp\left(-N\beta f[\bar{u}_{k}]\right)$ (5)
where $f[\bar{u}_{k}]=\min_{\left\{u_{k}\right\}}f[u_{k}]$, with $f[u_{k}]=f(u_{0},\ldots,u_{M-1})$ and $\beta=1/k_{\rm B}T$, with $k_{\rm B}$ Boltzman’s constant and $T$ the temperature, and
$A=\left({\left(\frac{\beta}{2}\right)^{M}\,{\det}H_{f}[\bar{u}_{k}]}\right)^{-1/2}\prod_{p=0}^{M-1}\sqrt{\frac{D_{p}}{\beta}}\;,$ (6)
with $H_{f}[\bar{u}_{k}]$ the Hessian of $f[u_{k}]$ at its global minimum. By noticing that in the thermodynamic limit $f[\bar{u}_{k}]$ is the free energy per particle: $f[\bar{u}_{k}]=\lim_{N\to\infty}-(N\beta)^{-1}\ln Z$, it becomes apparent that computing the partition function is reduced to a minimization of the variational free energy per particle
@@ III.2 Ground State Phase Diagram
In order to study the ground state of the model (See App. C for the finite temperature phase diagram), we define the variational ground-state energy as the zero-temperature limit of the variational free energy $e_{0}(u)=\lim_{\beta\to\infty}f(u)$, such that
@@
It can be shown that the global minimum, $\bar{u}$, is proportional to the staggered magnetization (See App. D)
@@
We have thus found the existence of a first-order QPT on a finite portion of the critical line. We emphasize that this is valid for all strong long-range models, $\alpha<1$, and that this behaviour is still present at finite temperatures (See App. C). This constitutes a qualitative difference with respect to the nearest-neighbor limit, $\alpha\to\infty$, of the model, in which the QPT is second-order along the full critical line except for the point where $\omega_{z}=0$, when the model is classical [36, 35]. This is sketched in Fig. 2(b) for comparison. It is worth noting that there is some evidence that the phase transition could be mixed order [53, 54] around the classical point in the nearest-neighbor limit, but not purely first order like we have just described for the strong long-range regime [55].
@@
For $\alpha=0$ it is possible to do a classical analysis of the model that predicts the same phase diagram that we just described, with distinct finite regions of first- and second-order phase transitions. See App. E for more details. This illustrates that a classical analysis can serve as an exploratory tool for spin models with all-to-all interactions. It offers an alternative perspective on the phase transition based on the magnetizations of each sublattice rather than the order parameter (the staggered magnetization in this case). In any case, it is not a replacement for an exact solution due to the fact that its application is limited to $\alpha=0$ and that it does not allow for the computation of susceptibilities.
@@ III.3 Analysis of correlations
The susceptibility is proportional to the Kubo correlator [56][Chap. 4], and hence a measure of correlations between spins. For a translation invariant model the susceptibility can be computed analytically [57][Chap. 6] (see details in App. D). At zero temperature it can be written as
@@ IV Numerical solution for arbitrary interaction range
Having established a significant difference between the phase diagram of the strong long-range and nearest-neighbors models, it is natural to wonder how this difference is interpolated for arbitrary values of the range of interactions, $\alpha$. Since only the strong long-range regime is analytically tractable, in this section we resort to zero-temperature variational quantum Monte Carlo (qMC) [61] simulations with a visual transformer (ViT) [38, 62, 63] ansatz for finite system sizes to study the model in the full range of interactions. We show the results obtained for spins with $s=1/2$ using this architecture, given its recent success in describing spin-1/2 chains with long-range interactions [30]. Details on the particular implementation used here can be found in App. G.
@@
In Fig. 4 we observe how $\langle m_{s}^{2}\rangle$ behaves for different slices of the phase diagram (cf. Fig. 2) as a function of the different values of $\alpha$ for a fixed chain size of $N=50$. In panels (a), (b) and (c), which correspond to vertical slices (fixed transverse field), we see that for certain values of $\alpha$ the nature of the phase transition evolves from first to second order as the transverse field, $\omega_{z}$, increases. Panel (a), corresponding to $\omega_{z}/(s\Gamma)=0.2$, merits special attention. It shows that the discontinuity characteristic of first-order phase transitions is clearly present up to $\alpha=2.5$ (See App. H for further confirmation). In panels (b) and (c) we see how this transition becomes second order, with a continuous change of the order parameter at the critical point. If instead of comparing a given value of $\alpha$ across panels we focus on any given panel, (a), (b) or (c), we observe that the transition develops a discontinuity as $\alpha$ is lowered. In all cases the transition is clearly discontinuous for $\alpha\leq 1$, in agreement with the analytical predictions.
@@
To better certify the order of the different phase transitions, beyond a visual analysis of the discontinuities (or lack thereof) of the order parameter, we perform a finite-size scaling analysis. In Fig. 5 we show this analysis for $\alpha=2$ (See App. H for other values of $\alpha$). As it can be seen from panel (a), all the curves corresponding to $N=50,70,100$ coincide, showing an independence with size characteristic of first-order transitions, in addition to the discontinuity in the order parameter. In contrast, in panel (c), the curves exhibit the typical size dependence of second-order phase transitions. The continuous change of the order parameter becomes increasingly non-analytical as the size increases. The crossover point marks the second-order critical point expected in the thermodynamic limit.
@@
Figure 6 summarizes the numerical results discussed in this section, along with additional results reserved for App. H. We sketch the critical line for different values of $\alpha$, highlighting the regions where it is first and second order. We have established that the first-order phase transition at finite transverse field is present for $\alpha\leq 2.5$. The portion of the critical line that is first order decreases progressively with increasing $\alpha$ until for $\alpha\gtrsim 3$ the full critical line is second order. This is remarkable because it indicates that the first-order phase transition is present not only in the strong long-range regime but also beyond the mean-field threshold and presumably (see the next paragraph) in the full weak long-range regime. Although the precise boundaries, $\alpha_{\rm MF}$ and $\alpha^{*}$, between the regime in which the model exhibits mean-field critical exponents and the weak long-range regime and between the weak long-range and short-range regimes, have not been established for the antiferromagnetic model under consideration, it is known that for the analogous ferromagnetic model they lie at $\alpha_{\rm MF}=5/3\approx 1.66$ and $\alpha^{*}=3$ [8]. Additionally, we have witnessed the shrinking of the antiferromagnetic phase as the model approaches the short range regime, driven by a decrease of the critical transverse field at zero longitudinal field from $\omega_{z}/(s\Gamma)=2$ for $\alpha>1$ to $\omega_{z}/(s\Gamma)=0.5$ for $\alpha\gtrsim 7$.
@@ V Discussion
The fact that similar behaviour has been observed in a model with only antiferromagnetic nearest-neighbor and ferromagnetic next-nearest-neighbor interactions [31] suggests that the key ingredient that stabilizes a first-order phase transition is the presence of intrasublattice ferromagnetic interactions, in contrast with models with only intersublattice antiferromagnetic interactions. To verify this hypothesis we have also considered Hamiltonian (1) but without intrasublattice ferromagnetic interactions, i.e. with $J_{ij}=0$ for $i+j$ even. This model is not tractable with the analytical method used in Sec. III, so an exact solution in the strong long-range regime is not possible. Nevertheless, a classical analysis of the $\alpha=0$ limit, analogous to the one described in App. E, shows that the phase transition is always second order, except for the point of vanishing transverse field, $\omega_{z}=0$, where the model is classical. Thus, we can conclude that intrasublattice ferromagnetic interactions (on top of intersublattice antiferromagnetic interactions) are the key ingredient to generate a first-order phase transition. In that sense, in the tunable-range unfrustrated antiferromagnetic Ising model of Hamiltonian (1), $\alpha$ acts as a tuning knob to change the ratio between inter- and intrasublattice interactions. For low enough $\alpha$, intrasublattice ferromagnetic interactions become large enough to stabilize a first-order phase transition. The same mechanism might explain the first-order phase transitions reported in other quantum long-range models with staggered interactions [20, 21, 22].
@@ Appendix A Solving the staggered antiferromagnetic Ising chain
$Z=\int\prod_{k=0}^{M-1}\sqrt{\frac{ND_{k}}{\pi\beta}}du_{k}\,\exp\left\{-N\beta f[u_{k}]\right\}\;,$ (19)
@@
Appendix C Analytical study of the non-zero temperature phase diagram
In the strong long-range regime, we have reduced computing the canonical partition function to solving a minimization problem. We have to find the global minimum of the variational free energy per particle, Eq. (7), with respect to the auxiliary fields $\left\{u_{k}\right\}$. We have discussed that the problem becomes univariate for every $\alpha<1$, i.e. in the global minimum $\bar{u}_{k}=u_{0}\delta_{k0}$. Although in the main text we focus on the ground state phase diagram, we can also use this technique to study the finite temperature properties of the model 32.
In the ground state phase diagram we have found the existence of a critical line whose order shifts from first to second order at a tricritical point. This is in contrast with the nearest-neighbors limit, where the critical line remains of second order except on the vanishing-transverse-field limit. In this appendix we will show that this long-range behavior of the phase diagram is still present for finite temperatures.
To study the phase diagram we numerically minimize the univariate free energy per particle, $f(u_{0})$. To find the second-order critical line and the tricritical point we also use Landau theory. In Fig. 7(a), (b) and (c) we show the phase diagram obtained via numerical minimization of the variational free energy and the second-order critical line obtained using Landau Theory.
Figure 7: (a) Comparison between the phase diagram obtained via numerical minimization of the variational free energy and the second order critical line obtained using Landau theory, using $\beta s\Gamma=200$. (b) $\beta s\Gamma=1.6$. (c) $\beta s\Gamma=1.1$. (d) Thermal evolution of the second-order critical line and the tricritical point.
In Fig. 7(d) we summarize the evolution of the second-order critical line as temperature increases ($\beta s\Gamma$ decreases). What we find is that increasing temperature reduces the ordered phase, and it also reduces the portion of the critical line corresponding to a first-order phase transition. Up to the numerically achieved accuracy, the first-order critical line disappears for $\beta s\Gamma<3/2$, and the ordered phase finally disappears for $\beta s\Gamma<1$.
Appendix D Analytical calculation of the magnetization and the susceptibility
@@
Appendix E Classical analysis of the phase transition for $\alpha=0$
@@
This Hamiltonian commutes with the total spin operators, $J_{\Lambda}^{2}=(J_{\Lambda}^{x})^{2}+(J_{\Lambda}^{y})^{2}+(J_{\Lambda}^{z})^{2}$. Consequently, it connects only states with the same total spins $J_{A}^{2}$ and $J_{B}^{2}$. Using exact diagonalization for finite sizes, we show in Appendix F that the ground state always lies on the subspace of maximum total spins. This implies that the ground state properties of the model are perfectly captured by a model of two spins of sizes, $j_{A}=j_{B}=sN/2$, interacting according to Eq. (40). In the thermodynamic limit, $N\to\infty$, the description is further simplified by the fact that these spins can be described exactly in their classical limit [69]. We can therefore study the energy resulting from replacing the spins with classical magnetizations $J_{\Lambda}^{\gamma}\to sN/2m_{\Lambda}^{\gamma}$. The magnetizations are unit vectors that we can assume to lie on the $x,z$ plane $\boldsymbol{m}_{\Lambda}=(\cos\theta_{\Lambda},0,\sin\theta_{\Lambda})$. With this, the energy per site reads
@@
Appendix F Verification using exact diagonalization
For $\alpha=0$ there are two equivalent descriptions of the model in terms of either individual spins at each site, as in Hamiltonian (1), or in terms of total spin operators of each sublattice, as in Hamiltonian (40). As explained in App. E, the total spin of each sublattice, $J_{\Lambda}^{2}$, is a conserved quantity and the Hamiltonian is block diagonal with blocks of fixed total spins. Here we diagonalize the subspace of maximum total spins, $j_{A}=j_{B}=sN/2$ and compare its low-energy spectrum to the low-energy spectrum of the full Hamiltonian (1). In Fig. 9 we show that the ground-state energies of the two Hamiltonians coincide, with differences appearing in the multiplicity of excited states. We show this for $N=10$ and $s=1/2$ fixing $\omega_{x}/(s\Gamma)=0.2$ and varying $\omega_{z}/(s\Gamma)$ and vice versa, but the same behavior is observed for any accessible system size, $N$, and spin size, $s$, and for any values of $\omega_{x}/(s\Gamma)$ and $\omega_{z}/(s\Gamma)$ of the phase diagram. This implies that the low energy physics are well captured by the subspace of maximum total spins, $j_{A}=j_{B}=sN/2$. This fact is exploited in App. E to provide a classical description of the ground state phase diagram.